# The Generic Temperature Response of Large Biochemical Networks

### v1

@@ III Discussion
Here we have shown that the generic temperature response of large biochemical networks is quadratic in the Arrhenius plot, if the network has an imbalance of activation energies in forward versus background directions, as it is to be expected for developmental systems. In order to make general statements on such networks, we resorted to a statistical description which does not require knowledge of all activation energies; rather we only make assumptions about their distribution, which is a reasonable description for a large system size $N$, where not all degrees of freedom can be known. Our discussion was based on a general expression of the MFPT for arbitrary networks, Eq. (19), which follows from graph theory [72] and here has been derived in a pedagogical manner. We then rewrote this formula in the language of statistical physics. The coefficients appearing in the Taylor-series of the logarithmic MFPT can be interpreted in terms of the cumulants of the distribution of the sums along the spanning trees and two-tree forests. If this sum can be sufficiently well approximated by a Gaussian distribution, this yields a quadratic dependence in the Arrhenius plot, as described by Eq. (38). This is our main result, namely the quadratic response as the natural description for complex biochemical networks. This result would not apply for a network which has similar distributions for the activation energies for forward and backward directions, but this is unlikely for developmental systems.
@@
Crapse et al.[26] suggested that the observed deviation from a pure Arrhenius-equation was due to the biochemical reactions not following an ideal Arrhenius-behavior, rather than a result of the complexity of the system. They reached that conclusion based on the simulation of a linear chain of $1.000$ reactions described by an Arrhenius-equation, which failed to reproduce the observed quadratic response in the Arrhenius plot. They also conducted an experiment measuring the conversion of NAD+ to NADH via GAPDH catalysis and found that this reaction does not follow a pure Arrhenius-like dependence. However, closer inspection of their simulations shows that the activation energies were chosen uniformly in an energy range from $20\frac{\text{kJ}}{\text{mol}}-100\frac{\text{kJ}}{\text{mol}}$ and the prefactors independently at $T_{0}=295\text{K}$. Due to the absence of backward rates and branching in this model, this leads to Eq. (37), for which the uniform distribution results in a series with infinitely many non-vanishing terms. If the activation energies were chosen from a normal distribution, this would yield a quadratic dependence also on this level. In contrast, our results are much more general and require less assumptions.
@@
An important aspect of biochemical systems is that above the melting temperature of the involved proteins, the denaturation of the involved enzymes slows down the reaction rates quite substantially, which is not described by the Arrhenius equation [80, 81]. This could explain the curvature of the Arrhenius plots seen at small values of the inverse temperature, in particular for temperatures of about $30^{\circ}\text{C}$ and above, where we observe that the high temperature data points (low inverse temperature) in Fig. 5-7 all lie below the quadratic curve. In the measurement by Powsner on the pupal development rates (Fig. 7B), the decrease is most notable, and it is not unlikely that this indicates that the denaturation threshold for a protein that is important for the developmental progression is surpassed at these temperatures. On the other hand, however, this observation also suggests that the larger part of the curve should not be affected by protein denaturation, and that our theory indeed captures the generic behaviour of this system. It should be mentioned that Powsner already noticed in his work that the combination of a few reactions, each described by an Arrhenius equation, does not necessarily yield a global Arrhenius equation and hypothesized that the observed temperature response may be a consequence of the complexity of the underlying biochemical network. However, he did not give an mathematical description as provided here. In the future, it might be interesting to extend our theory by the effect of protein denaturation, e.g. by introducing it as additional state in the network.
The observation that the development rates can be well described by a quadratic fit in the Arrhenius plot does not seem to be limited to Drosophila. Crapse et al.[26] provide similar results for the development of Xenopus laevis embryos, and Ratkowsky et al.[27] found empirically that the growth of bacterial colonies can be well described by a quadratic fit in an Arrhenius plot.
@@
The discussion in this work was limited to the mean of the first-passage time, although our formalism in principle covers all higher moments, too. The focus on the first moment seems justified by the observation that the coefficient of variation (standard deviation divided by mean) for the development times of Drosophila embryos is in the order of a few percent across different temperatures[78]. This results most likely from a strong forward bias for the rates of a developmental network with many steps like cell cycle checkpoints and cell division being practically irreversible. It is well known and follows from our formalisms that this tends to yield sharp FPT distributions around the mean in the limit of large networks[77] (compare also the example in appendix B).
@@
The framework developed in this work was motivated by describing the temperature dependence of complex networks where not all degrees of freedom are known. Here we focused on development, but as mentioned in the introduction, there exist other complex biological systems with interesting temperature dependence, in particular in the contexts of fever and climate change. We also note that our results are relatively general statements about the mean first passage time of master equations and are therefore not per se limited to biochemical networks and temperature effects. The idea to describe large reaction networks from a coarse-grained perspective with appropriate approximations has been discussed elsewhere [82], but neither the mean first passage time nor the possibility to describe the rate constants on a statistical basis the way proposed here seem to be explored so far. First passage-time problems are ubiquitous in biology and biochemistry [83] and it is very interesting how far one can get without knowing all the details of the system under consideration.
@@ Appendix A Obtaining the MFPT by counting
To illustrate the application of the graph theoretical approach used here, we derive the MFPTs for two examples which also can be obtained by applying Laplace-transforms[77].
@@
Note that there is only one spanning tree with all flow directed towards $N+1$, namely the one consisting only of the forward rates, meaning that:
@@
The $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ are the chains which start off towards the right and switch their direction at $j$ and possibly also at a later vertex $m$. This is shown in Fig. (9).
@@
Appendix B First-Passage Times for a Linear Chain of Reactions
The simplest example of a large network is a linear chain of irreversible reactions, i.e., $k_{i,j}=k_{i}\delta_{i+1,j}$, as illustrated in Fig. (12)
$\sum_{\mathcal{T}_{[N+1]}}w(\mathcal{T})=\prod_{i=1}^{N}k_{i}.$ (47)
Figure 13: For $i\leq j$, the only corresponding spanning forest is the one without the $k_{j}$-edge. For $i>j$, there is no such tree.
Then,
One finds for the first two moments:
$\langle\tau_{i}\rangle$ $=\sum_{j=i}^{N}\frac{1}{k_{j}},$ (50)
which yields the standard deviation:
$\sigma_{\tau_{1}}=\Big{(}\sum_{j=1}^{N}\frac{1}{k_{j}^{2}}\Big{)}^{\frac{1}{2}}$ (52)
and the coefficient of variation (CV) reads:
If $k_{i}=k$ $\forall i$, one finds in general:
$\langle\tau^{n}\rangle_{i}=\frac{1}{k^{n}}\frac{(N+n-i)!}{(N-i)!},$ (54)
which are precisely the moments for an $N-i+1$-dimensional Erlang distribution ($\Gamma$-distribution with non-negative integer as shape parameter) with $\lambda=\frac{1}{k}$:
$f^{\tau_{i}}(t)=\frac{k^{N-i+1}}{(N-i+1)!}t^{N-i}e^{-kt},$ (55)
as it should be since this is precisely the distribution that arises from summing up $N-i$ iid exponentially distributed random variables.
Note that the coefficient of variation (CV) reads:
Appendix C Fitting parameters for quadratic fits in main text
@@
The data in Fig 5-7 were fitted to a quadratic function of the shape $\ln{k}=aT^{-2}+bT^{-1}+c$. The corresponding parameters are listed in the two following tables:
Figure $a$ $[10^{7}K^{2}\ln{\text{min}^{-1}}]$ $b$ $[10^{5}K\ln{\text{min}^{-1}}]$ $c$ $[10^{2}\ln{\text{min}^{-1}}]$
5(a) -1.06 0.68 -1.10
5(b) -1.81 1.19 -1.98
5(c) -1.96 1.30 -2.17
5(d) -1.71 1.13 -1.88
Table 1: Fitting parameters for the quadratic fit of the data from[26] shown in Fig. 5 in the main text
Figure $a$ $[10^{7}K^{2}\ln{\text{h}^{-1}}]$ $b$ $[10^{5}K\ln{\text{h}^{-1}}]$ $c$ $[10^{2}\ln{\text{h}^{-1}}]$
6(a) -1.48 0.97 -1.60
6(b) -1.47 0.97 -1.59
6(a) -1.48 0.97 -1.60
7(a) -1.74 11.39 -1.88
7(b) -2.62 1.73 -2.89
7(c) -2.36 1.56 -2.59
7(d) -1.80 1.18 -1.96
7(e) -1.75 1.15 -1.91

### v2

@@ III Discussion
Here we have shown that the generic temperature response of large biochemical networks is quadratic in the Arrhenius plot if the network has an imbalance of activation energies in forward versus backward directions. This seems reasonable for developmental processes and is supported by evidence from in vitro studies with enzymes from Xenopus showing a substantial imbalance in the activation energies of the cyclin synthesis and degradation rates [48].
@@
In order to make general statements on such networks, we employed a statistical description which does not require knowledge of all activation energies. Instead, we only make assumptions about their distribution, which is a reasonable description for a large system size $N$, where not all degrees of freedom can be known. Our discussion was based on a general expression of the MFPT for arbitrary networks, Eq. (19), which follows from graph theory [74] and has been derived here in a pedagogical manner. We then rewrote this formula in the language of statistical physics. The coefficients appearing in the Taylor-series of the logarithmic MFPT can be interpreted in terms of the cumulants of the distribution of the sums along the spanning trees and two-tree forests. If this sum can be sufficiently well approximated by a Gaussian distribution, this yields a quadratic dependence in the Arrhenius plot, as described by Eq. (37). This is our main result, namely the quadratic response as the natural description for complex biochemical networks. This result need not apply for a network which has similar distributions for the activation energies for forward and backward directions, but this is unlikely for developmental systems.
Being the result of evolution, the assumption that the activation energies can be described as independent from each other needs to be treated with care. But since there are many different reactions involved in a process such as the development of an embryo, some of which are purely biochemical (like the phosphorylations occuring in cell cycle control), some of which are more biophysical in nature (like the growth of microtubules), it seems reasonable to assume that most activation energies are indeed independent of each, especially when contained in different modules of the overall network. Furthermore, generalizations of the central limit theorem exist that allow for a weak degree of dependence, e.g. correlations decaying with distance in the network (see for example [87]). In the future, it might be interesting to experimentally determine all activation energies in a network of interest and then to predict the resulting temperature dependence using the formalism presented here.
The theory was validated numerically by drawing spanning trees and two-tree spanning forests from a random graph on $N_{v}=1.000$ vertices. Individual activation energies were drawn from homogeneous distributions, also in order to avoid using a Gaussian distribution from the start. We then computed the histograms of their total activation energies, which approach a normal distribution as expected from the central limit theorem (Fig. 7). The global rate constant in an Arrhenius plot is then of quadratic shape, and this quadratic dependence becomes more pronounced for larger $N_{v}$, as predicted by the theory (Fig. 7).
@@
Crapse et al. [24] suggested that the observed deviation from a pure Arrhenius-equation was due to the biochemical reactions not following an ideal Arrhenius-behavior, rather than a result of the complexity of the system. They reached that conclusion based on the simulation of a linear chain of $1.000$ reactions described by an Arrhenius-equation, which failed to reproduce the observed quadratic response in the Arrhenius plot. They also conducted an experiment measuring the conversion of NAD+ to NADH via GAPDH catalysis and found that this reaction does not follow a pure Arrhenius-like dependence. However, closer inspection of their simulations shows that the activation energies were chosen uniformly in an energy range from $20\frac{\text{kJ}}{\text{mol}}-100\frac{\text{kJ}}{\text{mol}}$ and the prefactors independently at $T_{0}=295\text{K}$. Due to the absence of backward rates and branching in this model, this leads to a different series (see section II.8), for which the uniform distribution yields infinitely many non-vanishing terms. If the activation energies were chosen from a normal distribution, this would yield a quadratic dependence also on this level. In contrast, our results are more general and require fewer assumptions. Moreover, the measurement of the temperature dependence of the rate of the NAD+ to NADH reaction seems to be well described by a linear fit in the Arrhenius-plot up to at least $35^{\circ}\text{C}$. Indeed, for higher temperatures, the denaturation of the involved enzymes significantly decreases the reaction rate. The measurements of the development rates only reached about $33^{\circ}\text{C}$, though. Therefore, this can most likely not explain the quadratic temperature dependence observed over the entire range.
@@
The theory developed so far is of course no longer valid above the melting temperature of the proteins. Denaturation slows down the reaction rates quite substantially and this is not described by the Arrhenius equation [85, 88]. This could explain the curvature of the Arrhenius plots seen at small values of the inverse temperature, in particular for temperatures of about $30^{\circ}\text{C}$ and above, where we observe that the high temperature data points (low inverse temperature) in Fig. 9-11 all lie below the quadratic curve. In the measurement by Powsner on the pupal development rates (Fig. 11B), the decrease is most notable, and it is not unlikely that this indicates that the denaturation threshold for a protein that is important for the developmental progression is surpassed at these temperatures. On the other hand, however, this observation also suggests that the larger part of the curve should not be affected by protein denaturation, and that our theory indeed captures the generic behaviour of this system. Future work could extend our theory to include the effects of protein denaturation, for example, by introducing it as an additional state in the network.
It should be mentioned that Powsner already noticed in his work that the combination of a few reactions, each described by an Arrhenius equation, does not necessarily yield a global Arrhenius equation and hypothesized that the observed temperature response may be a consequence of the complexity of the underlying biochemical network. However, he did not give a mathematical description as provided here.
The observation that the development rates can be well described by a quadratic fit in the Arrhenius plot does not seem to be limited to fly and frog as discussed above. Rombouts et al. [48] use it to fit the timing of the early cell cycles of Caenorhabditis elegans, Caenorhabditis briggsae, Danio rerio, Xenopus tropicalis and Xenopus laevis, and Aquilanti et al. [89] found that a quadratic dependence in the Arrhenius plot is a good description for data on the respiration rate of Camellia Japonica leaves.
Our result is a statement about large biochemical networks. Due to the exponential nature of the Arrhenius equation, it is reasonable to assume that this effect dominates more subtle temperature dependencies, like the dependence of the diffusion constant on temperature. According to the Einstein relation, the diffusion constant $D$ is linear in $T$ and a change from $T=300K$ to $T=310K$ increases $D$ by a mere 3 %, while chemical rates often double or triple under a change of $\Delta T=10K$ [90]. Therefore, it seems justified to compare the theory to data on developmental processes, without claiming to capture every nuance of the effect of temperature on these complex systems. If needed, the effect of temperature on the diffusion constant can be included by decomposing binding rates into diffusion and reaction parts [91].
@@
The discussion in this work was limited to the mean of the first-passage time, although the formalism in principle covers all higher moments, too. The focus on the first moment seems justified by the observation that the coefficient of variation (standard deviation divided by mean) for the development times of Drosophila embryos is in the order of a few percent across different temperatures [81]. This results most likely from a strong forward bias for the rates of a developmental network with many steps like cell cycle checkpoints and cell division being practically irreversible. It is well known that this tends to yield sharp FPT distributions around the mean in the limit of large networks [80] (compare also the example in section II.8).
@@
The framework developed here was motivated by describing the temperature dependence of complex networks where not all degrees of freedom are known. In this work, we focused on development, but as mentioned in the introduction, there exist other complex biological systems with interesting temperature dependence, in particular in the contexts of fever and climate change. Our results are relatively general statements about the mean first-passage time of master equations and are therefore not per se limited to biochemical networks and temperature effects. The algorithm used to obtain the distributions of total activation energies along trees and two-tree forests also provides a method for calculating the MFPT for large networks. It may be worthwhile to explore how this compares to obtaining the MFPT from simulations of stochastic dynamics [92], and whether it can also be applied to study other network properties related to graph theory.
The idea to describe large reaction networks from a coarse-grained perspective with appropriate approximations has been discussed elsewhere [93], but neither the mean first-passage time nor the possibility of describing the rate constants on a statistical basis as proposed here seem to be explored so far. First-passage time problems are ubiquitous in biology and biochemistry [94, 95] and it is very interesting how far one can get without knowing all the details of the system under consideration.
@@ Appendix A Obtaining the MFPT by counting
To illustrate the application of the graph-theoretical approach used here, we derive the MFPTs for two examples which also can be obtained by applying Laplace-transforms [80].
@@
Note that there is only one spanning tree with all flow directed towards $N_{v}$, namely the one consisting only of the forward rates, meaning that:
@@
The $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ are the chains which start off towards the right and change their direction at $j$ and possibly also at a later vertex $m$. This is shown in Fig. 18.
Figure 18: Counting the spanning forests rooted at $N+1$ $\mathcal{F}_{[j,N+1]}^{1\rightarrow j}$ for the on-step master equation shown in Fig. (17).
@@
Appendix B Fitting parameters for the figures in the main text
The parameters for the fitted normal distributions for the distribution of the total activation energies of the spanning trees and spanning forests in Fig. 7 are:
Figure $N_{v}$ $E_{\mathcal{T}}$ $\frac{\text{kJ}}{\text{mol}}]$ $E_{\mathcal{F}}$ $\frac{\text{kJ}}{\text{mol}}]$
7(a,b) 2020 $953.74\pm 51.22$ $914.26\pm 58.23$
7(d,e) $100$ $4888.04\pm 179.66$ $4845.02\pm 183.05$
7(g,h) $1.000$ $52756.76\pm 607.74$ $52696.46\pm 618.11$
Table 1: Fitting parameters for the normal distribution in Fig. 7.
@@ Appendix B Fitting parameters for the figures in the main text
Quadratic function of the shape $\ln{k}=aT^{-2}+bT^{-1}+c$ were fitted to the data in Fig 9-11. The corresponding parameters and the standard errors are listed in the following table:
Figure $a$ $10^{7}K^{2}]$ $b$ $10^{5}K]$ $c$ $10^{2}\ln{\text{h}^{-1}}]$
9(a) -$\textbf{-1.06}\pm 0.31$ $\textbf{0.68}\pm 0.21$ -$\textbf{-1.10}\pm 0.36$
9(b) -$\textbf{-1.81}\pm 0.13$ $\textbf{1.19}\pm 0.09$ -$\textbf{-1.98}\pm 0.15$
9(c) -$\textbf{-1.96}\pm 0.14$ $\textbf{1.30}\pm 0.09$ -$\textbf{-2.17}\pm 0.16$
9(d) -$\textbf{-1.71}\pm 0.27$ $\textbf{1.13}\pm 0.19$ -$\textbf{-1.88}\pm 0.31$
10(a) -$\textbf{-1.48}\pm 0.06$ $\textbf{0.97}\pm 0.04$ -$\textbf{-1.60}\pm 0.07$
10(b) -$\textbf{-1.47}\pm 0.06$ $\textbf{0.97}\pm 0.04$ -$\textbf{-1.59}\pm 0.07$
11(a) -$\textbf{-1.74}\pm 0.07$ $\textbf{1.14}\pm 0.04$ -$\textbf{-1.88}\pm 0.08$
11(b) -$\textbf{-2.62}\pm 0.18$ $\textbf{1.73}\pm 0.12$ -$\textbf{-2.89}\pm 0.21$
11(c) -$\textbf{-2.36}\pm 0.22$ $\textbf{1.56}\pm 0.15$ -$\textbf{-2.59}\pm 0.25$
11(d) -$\textbf{-1.80}\pm 0.04$ $\textbf{1.18}\pm 0.03$ -$\textbf{-1.96}\pm 0.06$
11(e) -$\textbf{-1.75}\pm 0.07$ $\textbf{1.15}\pm 0.04$ -$\textbf{-1.91}\pm 0.08$